OpenStudy (anonymous):

What is the standard form of y=2x-3?

OpenStudy (rulnick):

-2x+y=-3

Directrix (directrix):

Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers.

Directrix (directrix):

2x + (-1) y = 3

OpenStudy (anonymous):

y=2x-3 2x - y = 3